Light is traveling from air to an ‘unknown material’. The angle of incidence is measured at 30.0o​ ​, and the angle of refraction is measured at 26.3o​ ​. What is the speed of light in this ‘unknown material’.
1
Expert's answer
2020-11-30T14:54:42-0500

Let's first find the index of refraction of the 'unknown material' from the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,"

here, "n_i=1.0" is the index of refraction of the incident medium (or the air), "n_r"is the index of refraction of the refractive medium (or the 'unknown material'), "\theta_i=30^{\circ}" is the angle of incidence, "\theta_r=26.3^{\circ}" is the angle of refraction.

Then, from this formula we can find "n_r":


"n_r=\dfrac{n_isin\theta_i}{sin\theta_r}=\dfrac{sin30^{\circ}}{sin26.3^{\circ}}=1.13"

Finally, we can find the speed of light in the ‘unknown material’ from the Snell's Law:


"\dfrac{v_r}{v_i}=\dfrac{n_i}{n_r},""v_r=\dfrac{n_iv_i}{n_r}=\dfrac{1.0\cdot 3.0\cdot10^8\ \dfrac{m}{s}}{1.13}=2.65\cdot10^8\ \dfrac{m}{s}."

Answer:

"v_r=2.65\cdot10^8\ \dfrac{m}{s}."